OpenStudy (anonymous):

f(x)=3x+1 and f^-1=x-1/3 help me :(

OpenStudy (anonymous):

to find the inverse, you'll want to invert the x and y varbiables. (f(x) = y on graphs usually) f(x) = y = 3x + 1 invert x and y x = 3y + 1 put it back in the form of y = f(x) y = (x - 1)/3 = f^(-1) (x)